From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Feb 23 2003 - 23:53:29 EST
------------------------------------------------------------ Date: Sun Feb 23 23:53:29 2003, Shift: 16:00-24:00 Supervisor: MM, with: Log book#14 pages 34 - 34 Shift Summary: ------------------------------------------------------------ 1.There was a long store during this shift, but background rates were to high to take data; TOFWrate>300000 during entire store. MCR reported that there was some problem that they had not figured out, they were trying to steer the beam, but whitout beeing able to get the backgrounds down signifigantly. In the beginning of the shift we were in the midle of the store from last shift, and data run8050 was taken. MRS TPCs were of in this run due to high backgrounds. 2.Only pedestal runs. 3.Powersupply 1458A tripped turning of the power to TPCs and Si tiles. This problem was fixed. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Sun Feb 23 2003 - 23:55:04 EST